# Adaptive behavior with stable synapses

### v1

@@ Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated in machine learning and reinforcement learning to synaptic plasticity, and, in general, to changes and optimization of network parameters. However, such rapid changes are not coherent with the timescales of synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical network reconfiguration. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could also be supported by input segregation and dendritic amplification, features extensively observed in biological networks. We propose an architecture composed of gain-modulated recurrent networks that excels at in-context learning, showing abilities inaccessible to standard networks. We argue that such a framework can describe the psychometry of context-dependent tasks on humans and other species, solving the incoherence of plasticity timescales. When the context is changed, the network is dynamically reconfigured, and the predicted output undergoes dynamic updates until it aligns with the information embedded in the context. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
Introduction
@@
Despite their success, transformer architectures present many shortcomings, primarily in memory requirements, that scale quadratically with sequence length. This limits the scalability of transformers and hinders their applicability to tasks requiring the processing of long sequences, such as full-length document analysis or video understanding. To address these problems, several efficient linearized attention models have emerged, characterized by a forward pass executed in an RNN-like manner with constant inference memory costs. Recently, deep linear RNN architectures [12, 13] have yielded notable performance improvements over transformers, particularly in long-sequence tasks. Zucchet et al. [14] explore the efficacy of these deep linear RNNs, incorporating element-wise multiplications, in approximating attention mechanisms and facilitating in-context learning.
@@
Our investigation extends this premise, aiming to unravel how in-context learning, as observed in deep learning artificial neural networks, might also manifest in biological recurrent neural networks. Are there unique features in transformers that are also present in biological networks? We claim that in-context learning can be supported by input segregation and dendritic amplification, features extensively observed in biological networks. We argue that those are a biologically plausible ingredient capable of implementing a process that is similar to the attention mechanism present in transformers.
Recent findings on dendritic computational properties [15] and on the complexity of pyramidal neurons dynamics [16] motivated the study of multi-compartment neuron models in the development of new biologically plausible learning rules [17, 18, 19, 20]. It has has been proposed that segregation of dendritic input [18] (i. e., neurons receive sensory information and higher-order feedback in segregated compartments) and generation of high-frequency bursts of spikes [20] would support backpropagation in biological neurons. In [21] authors suggest that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks. They show a possible implementation of this in a two-level network, where the high-network produces the contextual signal for the low-network.
Here, we propose an architecture composed of gain-modulated recurrent networks that demonstrate remarkable in-context learning capabilities, which we refer to as ’dynamical adaptation’. Specifically, we illustrate that our biologically plausible architecture can dynamically adapt its behavior in response to feedback from the environment without altering its synaptic weights. We present results for supervised learning of stationary and temporal datasets, as well as for reinforcement learning.
This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behavior in both natural and artificial intelligence domains.
Methods
Dynamical adaptation
@@ Methods
$\begin{cases}\tau_{e}\,\dot{e}&=E(e,y,f)\\ \tau_{y}\,\dot{y}&=Y(y,w,x)\\ \tau_{w}\,\dot{w}&=W(w,x,e)\end{cases}$ (1)
@@
Dynamical adaptation for linear regression
Consider now the task of learning a linear relationship between input $x$ and output $y$. In the conventional approach, a neural network is initialized with random parameters, essentially making random guesses. The model estimation can be improved through iterative adjustments of the network parameters (e.g. the synaptic weights), following the gradient of errors evaluated on a training dataset $(x_{i}^{data},y_{i}^{data})$. One step of the gradient would update the weights of a linear model $y=wx$, following $\delta w=-\sum_{i}(y-y_{i}^{data})x^{data}_{i}$, where $\eta$ is the learning rate. By the end of the training process, the network aims at accurately predict the output $y^{test}$ for a new input $x^{test}$. An equivalent error reduction can be achieved by updating the target output $y_{i}^{data}$ to $y_{i}^{data}-wx_{i}^{data}=y_{i}^{data}-y_{i}^{train}$ (let’s denote this the latter quantity as $e_{i}$), where $y_{i}^{train}=wx_{i}^{data}$ is the current model estimation of $y_{i}^{data}$. $e_{i}$ can then be updated iteratively, and when it reaches the value zero, it indicates that the interpretation of the dataset is coherent, and the network predictions on new data $x^{test}$ are reliable. We can interpret this dynamics as changing the way $x^{test}$ is processed in function of the virtual weights $w$. Typically, when a network dynamically alters the way it receives an input, it is referred to as an attention mechanism. Following this concept, our goal is to project the dataset $(x_{i}^{data},y_{i}^{data})$ onto the network, influencing the way a new input $x^{test}$ (projected simultaneously) is processed. The information that drives and influences the attention mechanisms, $(x_{i}^{data},y_{i}^{data})$ in this case, is usually referred to as a context. When the input vector is one dimensional, the task can be solved by a network composed of three modules, one evaluating the error $e$, one evaluating the necessary update $\delta w$, and one evaluating the response $y$ to input data (see Fig. 1A). Their dynamics can be explicitly written as follows  [^1] Here we model the $y$ to be an instantaneous function of the variables $w,x$. This is equivalent to considering $Y(y,w,x)=xw-y$ in Equation (1) and taking the limit $\tau_{y}\to 0$. Similar considerations hold for $e$.:
$\begin{cases}e_{i}&=(y^{data}_{i}-x_{i}^{data}w)\\ \tau_{w}\,\dot{w}&=\sum_{i=1}^{M}x_{i}^{data}\,e_{i}\\ y&=x^{test}w\end{cases}$ (2)
with $e_{i}(0)=y_{i}^{data}-w_{0}x_{i}^{data}$, and $w_{0}$ are the initial weights. The conditioning dataset is $(x_{i}^{data},y_{i}^{data})$, with $i=1,...M$, while $(x^{test},y^{test})$ is one element of the test set. In this case one example ($M=1$) is sufficient to determine the response to new inputs. This formulation is equivalent to what derived in [11], with the difference that here many iteration are recursively evaluated over time, allowing the convergence to the gradient descent solution. See Results section for examples of applications. This approach can be extended to multidimensional linear datasets (see Supporting Information).
Dynamical supervised learning for temporal trajectory
The above formulation is well-posed for stationary dataset, however such a framework (the temporal convergence of the error to zero and of the response to the desired value) is tricky for temporal trajectories, which requires a little change of paradigm, to achieve a formulation that holds at each time step.
First of all we introduce a target temporal trajectory $y^{targ}(t)$, and define $y(t)=\bm{w}\cdot\bm{x(t)}$ as the current estimation of $y^{targ}(t)$, where $\bm{w}$ and $\bm{x(t)}$ are virtual weights and input vectors respectively, since we extended to formulation to a multidimensional dataset. In this case, there is not a stationary projection of the training set, but the current value of the target signal itself is projected at every time stem $t$. We do not separate the target estimation on a test set and on a training set, as a consequence, $y(t)=y^{train}(t)$.
This together with Eq. (2) allows us to write:
@@
In particular, we used two RNNs (see details below in the following Method subsections) that receive $\bm{x},\bm{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm).
In addition, we introduce a novel concept known as a gain-modulated network. This concept is inspired by the remarkable ability observed in biological neurons, particularly L5 neurons as discussed in previous studies (e.g., [22, 21]). These neurons exhibit the capacity to non-linearly integrate segregated inputs, a process critical for various cognitive functions. More details in the section "Gain Modulated Reservoir Computing". We empirically demonstrate that this network architecture outperforms standard RNNs in approximating and generalising the required virtual update rules described above, suggesting that gain modulation might be an important requirement for the adaptive behaviour observed in biological agents.
The formulation shown above can only be used to tackle linear problems, by learning linear relationships between $\bm{x}$ and $y$. However, if we consider that $\bm{x}$ is the activity of another RNN that operates as a reservoir computer extracting nonlinear features of the input sequence, a wider class of tasks can be tackled without changing the constraint of a linear readout $y=wx$.
Dynamical reinforcement learning
In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behaviour across diverse environments.
We start by defining a policy network, denoted as $\bm{\pi}=\mathrm{softmax}(\bm{y})$ which implements a policy mapping the agent state encoded by the vector $\bm{x}$ to a probability distribution over actions. For the sake of simplicity, We assume a linear agent such that $\bm{y}=\bm{w}\cdot\bm{x}$. This assumption is done without loss of generality since this could be easily extended by resorting to a reservoir computer as an intermediate layer as described above. The policy depends on the virtual weights $\bm{w}$, determined by the activity of an additional RNN. This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behaviour in real time. This can be formalized in a formulation that is very similar to the one used above, by changing the definition of $e(t)$ as follows:
@@
This formulation holds for a null discount factor [23, 24]). See Supplementary information for a complete derivation and for the description of the general case. The above equation can be rewritten in order to refer only to two networks, one for estimating the gradients, and one for the scalar product.
$\begin{cases}\tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq W(r,a,\bm{\pi})=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}\right)\odot\bm{x}\\ \bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\end{cases}$ (6)
Algorithm Distillation
@@
Gain modulated reservoir computing (GM-RC)
@@
where $\Theta$ represents trainable parameters. In the rest of the article, we will write $\bm{y}=\mathrm{RNN}_{\Theta}(\{\bm{x}\})$, dropping the temporal dependencies. 333 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [27].
@@
$\tau_{z}\dot{\bm{z}}$ $=\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(J\bm{z}+R\bm{x})\right)-\bm{z}$ (9)
@@
Where $\bm{b}^{\text{ap}}$ is a constant bias vector. Similarly, as before, the expression $\bm{y}=\mathrm{GMRNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir. As before, removing time dependencies we will have an instantaneous function $\bm{y}=\mathrm{GMNN}_{\Theta}(\bm{x}|\gamma\bm{x}^{\text{ap}})=\Theta\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(R\bm{x})\right)$ of the inputs $\bm{x}^{\text{ap}}$ and $\bm{x}$. We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the effect of the apical inputs on the network. Specifically, when $\gamma=0$, we obtain a network in which $\bm{x}^{\text{ap}}$ does not affect the gain modulation. The choice for the name $\bm{x}^{\text{ap}}$ is inspired by to current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high level information [22].
Results
Figure 1: Dynamical adaptation to changing context: A. Network architecture composed of 3 modules: $e$, $\delta w$ and $y$. The estimation on new data is produced by $y$, and is modulated by the error variable $e$ through $\delta w$. B. Example of the time course for the three variables in a single realization of the task. The values of $y$ converge to the correct values of (black dashed lines). C. This model accounts for the recovery of optimal performances (bottom, observed in humans and monkeys data) when the context suddenly changes, compared to a slower increase of performances achieved through synaptic plasticity.
Dynamical adaptation replaces synaptic plasticity
To demonstrate that suited responses can be provided relying on dynamical adaptation, we first consider a task requesting to learn a linear relationship between input $x^{data}$ and output $y^{data}$. As described in the Methods we consider a three module model (see Fig. 1A) described by the variables, $e$, $\delta w$, and $y$, whose dynamics is defined by Eq. (2), that is equivalent to following a gradient dynamics as respect to the parameters $w_{0}$. This update can be realized by treating $e$ as an input that, conditioned by the dataset $(x^{data},y^{data})$, evolves over time.
When $e\to 0$ the interpretation of the dataset is coherent, and the input-output mapping performed by the network on the new data is reliable. In this case one example is sufficient to determine the response to new inputs.
Fig. 1B shows the dynamics of three variables: the absolute value of $e$, and $\delta w$ as a consequence the error, are reduced over time, while the $y$ converges to the correct $y^{test}$ given the input $x^{test}$ (shown for 10 different new values).
We show the performances of the network whose parameters $w_{0}$ are trained through a gradient based plasticity on the rule $y^{data}=x^{data}$ (see first bar in the plots Fig. 1C), after the rule is switched to $y^{data}=-x^{data}$ (see, the second bar in which performances drop in both panels). In the top panel $w_{0}$ parameters are trained through gradient updates (we refer to Github repository https://github.com/cristianocapone/ABSS for details on parameters and implementation). In the bottom panel weights are not re-trained, but the last example (that is coherent with the new rule) is projected as a context (e.g. $(x^{data},y^{data})=(-1,1)$, starting the trial numbered as 1). This represents the agent understanding that the rule has been switched, and allows rapid and dynamic network reconfiguration and recovery of performances to optimal values, similarly to what observed in monkeys and humans data. An interesting example is the Wisconsin Card Sorting Test, in which participants are required to sort a deck of cards based on changing rules [30]. Initially, participants are given a rule (e.g., choosing by color) , and after a certain number of trials, the rule is changed without warning (e.g., choosing by shape). Participants then need to adapt to the new rule and adjust their strategy accordingly. The quick adaptation observed in this setup [31] reflects the ability to update mental models of the task, and flexibly apply new rules or strategies (similarly to Fig. 1C-bottom). On the other hand in some pathological subject, the adaptation after rule change appears to be slower [31], suggesting a learning process compatible with gradient-based plasticity (Fig. 1C-top).

### v2

@@ Abstract
Behavioral changes in animals and humans, as a consequence of an error or a verbal instruction, can be extremely rapid. Improvement in behavioral performances are usually associated in machine learning and reinforcement learning to synaptic plasticity, and, in general, to changes and optimization of network parameters. However, such rapid changes are not coherent with the timescales of synaptic plasticity, suggesting that the mechanism responsible for that could be a dynamical network reconfiguration. In the last few years, similar capabilities have been observed in transformers, foundational architecture in the field of machine learning that are widely used in applications such as natural language and image processing. Transformers are capable of in-context learning, the ability to adapt and acquire new information dynamically within the context of the task or environment they are currently engaged in, without the need for significant changes to their underlying parameters. Building upon the notion of something unique within transformers enabling the emergence of this property, we claim that it could be supported by gain-modulation, feature extensively observed in biological networks. We propose an architecture composed of gain-modulated recurrent networks that excels at in-context learning, showing abilities inaccessible to standard networks. We demonstrate that we can extend our approach to non-linear and temporal tasks and to reinforcement learning. Our framework contributes to understanding the principles underlying in-context learning and adaptive behavior in both natural and artificial intelligence.
1 Introduction
@@
Despite their success, transformer architectures present many shortcomings, primarily in memory requirements, that scale quadratically with sequence length. This limits the scalability of transformers and hinders their applicability to tasks requiring the processing of long sequences, such as full-length document analysis or video understanding. To address these problems, several efficient linearized attention models have emerged, characterized by a forward pass executed in an RNN-like manner with constant inference memory costs. Recently, deep linear RNN architectures [12, 13] have yielded notable performance improvements over transformers, particularly in long-sequence tasks. Zucchet et al. [14] explore the efficacy of these deep linear gated RNNs, incorporating element-wise multiplications, in approximating attention mechanisms and implementing in-context supervised learning.
@@
Our investigation extends this premise, aiming to unravel how in-context learning, as observed in deep learning artificial neural networks, might also manifest in biological recurrent neural networks. Are there unique features in transformers that are also present in biological networks? We claim that in-context learning can be supported by input segregation and dendritic amplification, features extensively observed in biological networks. We argue that those are a biologically plausible ingredient capable of implementing a process that is similar to the attention mechanism present in transformers. Recent findings on dendritic computational properties [15] and on the complexity of pyramidal neurons dynamics [16] motivated the study of multi-compartment neuron models in the development of new biologically plausible learning rules [17, 18, 19, 20]. It has has been proposed that segregation of dendritic input [18] (i. e., neurons receive sensory information and higher-order feedback in segregated compartments) and generation of high-frequency bursts of spikes [20] would support backpropagation in biological neurons. In [21] authors suggest that this neuronal architecture naturally allows for orchestrating “hierarchical imitation learning”, enabling the decomposition of challenging long-horizon decision-making tasks into simpler subtasks. They show a possible implementation of this in a two-level network, where the high-network produces the contextual signal for the low-network. Here, we propose an architecture composed of gain-modulated recurrent networks that demonstrate remarkable in-context learning capabilities, which we refer to as ’dynamical adaptation’. Specifically, we illustrate that our biologically plausible architecture can dynamically adapt its behavior in response to feedback from the environment without altering its synaptic weights. We present results for supervised learning of temporal trajectories and reinforcement learning, involving non trivial input-output temporal relations. This novel architecture aims to bridge the gap between biological-inspired in-context learning and the capabilities of artificial neural networks, offering a promising avenue for advancing our understanding of adaptive behaviour in both natural and artificial intelligence domains.
Our work generalizes and provides a biologically plausible implementation for the type of networks presented in [14]. Notably, our architecture has the same order of magnitude of trainable parameters and hidden units (see Appendix B). In our approach, in-context learning (ICL) can emerge simply by tuning the readout weights of the two involved networks. In contrast, previous implementations require biologically implausible mechanisms of temporal credit assignment for the network weights to converge correctly. We show that we can side-step this problem by dividing the architecture into two separate components and introducing an additional objective, which forces the first network to approximate the gradient of the second.
2 Methods
2.1 Dynamical adaptation replaces synaptic plasticity
@@
$\begin{cases}\tau_{e}\,\dot{e}&=E(e,y,f),\\ \tau_{y}\,\dot{y}&=Y(y,w,x),\\ \tau_{w}\,\dot{w}&=W(w,x,e).\end{cases}$ (1)
@@
2.2 Dynamical supervised learning for temporal trajectory
We consider the task of learning a target temporal trajectory $y^{targ}(t)$, We define $y(t)=\bm{w}\cdot\bm{x(t)}$ as the current estimation of $y^{targ}(t)$, where $\bm{w}$ and $\bm{x(t)}$ are virtual weights and input vectors respectively, since we extended to formulation to a multidimensional dataset. In this case, there is not a stationary projection of the training set, but the current value of the target signal itself is projected at every time stem $t$. We do not separate the target estimation on a test set and on a training set, as a consequence, $y(t)=y^{train}(t)$. In this case, learning can be formulated as follows:
@@ 2.2 Dynamical supervised learning for temporal trajectory
In particular, we used two RNNs (see details below in the following Method subsections) that receive $\bm{x},\bm{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm). In addition, we introduce a novel concept known as a gain-modulated network. This concept is inspired by the remarkable ability observed in biological neurons, particularly L5 neurons as discussed in previous studies (e.g., [22, 21]). These neurons exhibit the capacity to non-linearly integrate segregated inputs, a process critical for various cognitive functions. More details in the section "Gain Modulated Reservoir Computing". We empirically demonstrate that this network architecture outperforms standard RNNs in approximating and generalising the required virtual update rules described above, suggesting that gain modulation might be an important requirement for the adaptive behaviour observed in biological agents. The formulation shown above can only be used to tackle linear problems, by learning linear relationships between $\bm{x}$ and $y$. However, if we consider that $\bm{x}$ is the activity of another RNN that operates as a reservoir computer extracting nonlinear features of the input sequence, a wider class of tasks can be tackled without changing the constraint of a linear readout $y=wx$.
2.3 Dynamical reinforcement learning
In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behaviour across diverse environments. We start by defining a policy network, denoted as $\bm{\pi}=\mathrm{softmax}(\bm{y})$ which implements a policy mapping the agent state encoded by the vector $\bm{x}$ to a probability distribution over actions. For the sake of simplicity, We assume a linear agent such that $\bm{y}=\bm{w}\cdot\bm{x}$. This assumption is done without loss of generality since this could be easily extended by resorting to a reservoir computer as an intermediate layer as described above. The policy depends on the virtual weights $\bm{w}$, determined by the activity of an additional RNN. This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behaviour in real time. This can be formalized in a formulation that is very similar to the one used above, by changing the definition of $e(t)$ as follows:
@@ 2.3 Dynamical reinforcement learning
This formulation holds for a null discount factor [23, 24]), we refer to the Appendix section C for the description of the general case. The above equation can be rewritten in order to refer only to two networks, one for estimating the gradients, and one for the scalar product.
$\begin{cases}\tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(r,a,\bm{\pi})\simeq W(r,a,\bm{\pi})=r\left(\bm{\mathds{1}}_{a}-\bm{\pi}(\bm{y})\right)\odot\bm{x}\\ \bm{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\bm{x},\bm{w})\simeq Y(\bm{x},\bm{w})=\bm{x}\cdot\bm{w}\end{cases}$ (5)
2.4 Algorithm Distillation
@@
2.5 Gain modulated reservoir computing (GM-RC)
@@
where $\Theta$ represents trainable parameters. In the rest of the article, we will write $\bm{y}=\mathrm{RNN}_{\Theta}(\{\bm{x}\})$, dropping the temporal dependencies. 222 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [27].
@@
$\tau_{z}\dot{\bm{z}}$ $=\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(J\bm{z}+\beta\cdot R^{\text{ap}}\bm{x}^{\text{ap}}+R\bm{x})\right)-\bm{z}$ (8)
@@
Where $\bm{b}^{\text{ap}}$ is a constant bias vector. The hyperparameters $\alpha,\beta,\gamma$ modulate the effect of the gain modulation of the apical inputs. Specifically, when $\gamma=0$, we obtain a network in which $\bm{x}^{\text{ap}}$ does not affect the gain modulation. Similarly, as before, the expression $\bm{y}=\mathrm{GMRNN}_{\Theta}(\{\bm{x}\}|\{\gamma\bm{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir. As before, removing time dependencies, we will have an instantaneous function $\bm{y}=\mathrm{GMNN}_{\Theta}(\bm{x}|\gamma\bm{x}^{\text{ap}})=\Theta\phi\left(\left(\alpha\bm{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\bm{x}^{\text{ap}}\right)\odot(R\bm{x})\right)$ of the inputs $\bm{x}^{\text{ap}}$ and $\bm{x}$. We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the gain modulation effect of the apical inputs on the network. The choice for the name $\bm{x}^{\text{ap}}$ is inspired by the current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high-level information [22].
3 Results
Computing resources
The experiments of this paper were executed on a Macbook pro M3 CPU 12-core with 36 GB of RAM, and on a Macbook pro 2,9 GHz 6-Core Intel Core i9 with 32 GB of RAM.